A disc jockey must play 14 commercial spots during 1 hour of a radio show. Each commercial is either 30 seconds or 60 seconds long. If the total commercial time during 1 hours 11 minutes, how many 30-second commericals were played that hours? How many 60-second commercials?
Answer by stanbon(75887)

A disc jockey must play 14 commercial spots during 1 hour of a radio show. Each commercial is either 30 seconds or 60 seconds long. If the total commercial time during 1 hours 11 minutes, how many 30-second commericals were played that hours? How many 60-second commercials?
-----------------
Let # of 30 second spots be "t":
Let # of 60 second spots be "s":
----------------
EQUATIONS:
Quantity equation: t + s = 14
Time equations : 30t + 60s = 71 seconds
---------------------
t = 14-s
Substitute to get:
30(14-s)+ s = 14
420-29s = 14
29s = 406
s = 14 (# of sixty second spots)
----------------
Substitute into t = 14-s to get:
t = 14-14 = 0 (# of thirty second spots)
=============
Cheers,
Stan H.